# Sheshathanks: Shesha or Sheshanaga is a serpentine demigod in Hindu mythology with multiple heads and serves as the celestial bed of Lord Vishnu.  : Multi-head Microarchitectural Leakage Discovery in new-generation Intel Processors

### v1

@@ 4 Uncovered Transient Execution Paths
We ran a test campaign of $24$ hours, in which most observations were obtained within the first $8$ hours (refer to Algo. E in Appendix for description of the algorithm). Since we rely upon Intel Performance Monitoring Unit (PMU), which captures the events occurring in the system with certainty, the false positive rate of Shesha is zero. In this section, we detail the novel transient execution paths uncovered by Shesha as a result of the testing campaign. We also reverse engineer the root-cause of these transient execution paths and detail their effect on the overall behaviour of the system under test.
@@ 5 Shesha Design Choices
We now explain the choice of parameters by comparing the role of Particle behaviour and Swarm behaviour in Shesha. To do so, we define $6$ variants of Shesha wrt. hyperparameters $\beta$ (which controls Particle behaviour) and $\gamma$ (which controls Swarm behaviour). To compare between the variants, we define two metrics: ① time taken to find the first SIMD-Vector speculation, and ② time taken to find the first precision intermixing speculation. The reason for this (as detailed in subsequent sections) is because SIMD-Vector speculation is the easiest of the four (cf. Tab. 2) to discover. This is because of the large number of instructions under these ISEs, as well as no special requirement (like read-after-write dependencies, or specific floating-point arithmetic) needed to trigger this speculation. Precision intermixing, however, is on the other extreme, and is the hardest of the four to uncover. This is because (as we detail in Sec. 6.2) such speculation requires closely tied instructions that handle single/double precision respectively, and have read-after-write dependency.
Variants Shesha{β=1,γ=0} and Shesha{β=0,γ=1} are essentially equivalent to random sampling, with different sources. The former prioritizes mutating the chosen instruction within the particle itself by replacing it with a random instruction with probability $1$. The latter also works to similar goal, but replaces the chosen instruction with another instruction from other particle in the swarm. Then, variants Shesha{β=0.4,γ=0} and Shesha{β=0.1,γ=0} test the extent of particle behaviour by changing the probability $\beta$ of replacing a chosen instruction with a randomly sampled instruction. Finally, variants Shesha{β=0.1,γ=0.1} and Shesha{β=0.1,γ=0.4} test the extent of swarm behaviour by changing the probability $\gamma$ of replacing a chosen instruction with an instruction from another particle in the swarm. We summarize these in Tab. 3.
@@
As a result, in our final deployment of Shesha, we adopt a mixture of variant Shesha{β=0.4,γ=0} and variant Shesha{β=0.1,γ=0.4}. The first variant forms the Cognitive phase (cf. Sec. 3.4.2), while the second variant forms the Mixed phase (cf. Sec. 3.4.3). Finally, we note that variant Shesha{β=1,γ=0} simulates other fuzzing (essentially relying upon a random sampling strategy) based tools like [56, 39, 37, 40] had they supported ISEs like AVX/SSE/FMA and so on. For comparison however, we report the raw data from these works, albeit with different objectives over mostly x86 ISA. For instance, wrt. the speculative paths and ISA explored by [39], it is able to detect violations in about $5$ minutes (which is comparable to Shesha{β=0.4,γ=0}’s detection time for SIMD-Vector transient path detection). Likewise, in [37], the entire fuzzing campaign lasts about $20+$ CPU hours. On similar lines, the detection in [40] ranges from a minute to about an hour based on the type of speculation targeted 111[56] does not report convergence time.. One must note that the objectives of each of these tools (as presented in their corresponding papers) were different, and thus, they explore different subsets of ISAs. Shesha works with comparatively larger ISA set and thus explores wider range of instruction combinations and stimulates variety of specialized execution units and buffers.
@@ 6.1 SIMD-Vector Transition based SMC
SIMD and Vector instruction extensions like AVX, AVX2, AVX512EVEX, AVX512VEX, SSE, SSE2, SSE3, SSE4, SSE4a, and SSSE3 are central to several performance gains experienced in modern Intel processors. These ISEs allow greater than $64$ bit registers holding more than one data point, to be operated in parallel by the same instruction (hence the name SIMD type ISEs). To do so, modern IA-64 architectures define a register bank with sixteen $512$-bit wide registers. Interestingly, ISEs like AVX2/SSE allow differing representations of the same hardware register in two dimensions: ① width $w$ of the register, and ② the SIMD representation of the data contained in that width $w$. To elaborate on dimension ①, the same hardware registers can be interpreted as $512$-bit wide registers zmm0, zmm1, zmm2, ..., zmm15 (by ISEs like AVX512EVEX, AVX512VEX), as $256$-bit wide registers ymm0, ymm1, ymm2, ..., ymm15 (by ISEs like AVX2), and as $128$-bit wide regisers xmm0, xmm1, xmm2, ..., xmm15. Likewise, regarding dimension ②, the same data in any zmm/ymm/xmm register can be interpreted as differently packed integer/floating-point value. This discussion raises a concerning question: Given a processor allowing executions of ISEs requiring different representations of the same hardware, how are transitions between such ISEs handled? According to the Intel Software Developer Manual [23], the exact implementation to handle such transitions depends upon the exact micro-architecture under question. We defer the details of these implementations and our reverse-engineering approach to Appendix B. We make the following observation as the root cause of bad speculation (as in Tab. 2) encountered in modern Intel processors when transitioning from SIMD to Vector instructions (and vice-versa).
@@ 6.2 Precision based Transient Execution Paths
Precisely, for the discussed transient paths to kick in, the requirement is a set of at least two instructions with a Read-after-Write (RAW) dependency connecting them (using the xmm0 register in the above examples), without occurrence of an explicit conversion instruction (like CVTPD2PS or CVTPS2PD) in between. We note here that Write-after-Write (WAW) or Write-after-Read (WAR) dependencies do not trigger any of the following transient paths, simply because there is no need for any implicit precision conversion as the second instruction is writing to the dependent register and not reading from it. Writes from a single-precision instruction to a register holding a double-precision value are architecturally simply overwriting the registers, and thus incur no transient executions. We defer the detailed discussion on each of these transient execution paths and our reverse-engineering approach to Appendix A. We make the following observation as the root cause of all the bad speculation events when transitioning between single and double precision.
@@ 7.1 Transient Execution Window
Transient execution paths guarantee flushing of the in-flight micro-operations from the pipeline and re-steer the execution on the correct path. As such, the width of the transient window for any given transient execution path determines the number of micro-operations an attacker can issue on the transient path before the rollback of the instructions. Practically, a larger transient window allows for executing complex attack vectors in the transient path. Architecturally, no effect of any transient micro-operation is visible. Therefore, we need to rely upon micro-architectural effects to measure the transient window. In our case, we compute the transient window by counting the number of transient loads executed before the pipeline rollback. When no bad speculation (discovered by Shesha) occurs, we observe no activity in the transient path, implying absence of misprediction-based transient execution. Our results are summarized in Tab. 2. In comparison to the machine clears presented in [41], we conclude that the SIMD-Vector transition-based SMC exhibits a larger transient window than all other variants of transient executions through machine clears or assists (both uncovered here as well as in [41]). This observation directly ties in with the number of repetitions (for warming up the execution pipeline) an attacker needs to do on the transient path before the actual attack can be mounted (cf. Sec. 7.1). As evident, due to the largest transient window with minimum number of repetitions, SIMD-Vector transition-based SMC provides a very practical transient execution path (with effectively no repetitions required), thereby becoming an actual choice for an adversary to mount in-the-wild attacks.
Enlarging the transient window. As different transient execution paths have differing transient windows, not all of them would be directly suited to leak reliably. To circumvent this, a usual method is to warm up the pipeline before the actual attack; this can be done solely on the attacker’s end. This can be done by several repetitions (rep) of the same transient operations to maximize the transient window. To determine the extent of pipeline warm-up needed to leak data, we increase the repetitions of warm-up until we observe leakage from the Flush+Reload [58] covert channel. In Tab. 4, we provide a comparative analysis of the number of repetitions needed to warm up the pipeline. Lesser the number of required repetitions rep, more likely it is to find in-the-wild occurrence of the uncovered transient paths in in-the-wild code bases  [^2] We could not test the pipeline warmup and leakage rates for  [56] and [45] because of absence of Intel TSX on our test platforms..
@@ 7.2 SMT Covert Channel
We demonstrate a covert channel using contention in the affected hardware assets. Two processes are pinned to two logical cores of the same physical core. Hence, we assume Simultaneous Multi-Threading (SMT) or HyperThreading to be enabled for the covert channel construction to work since the affected assets identified (in Tab. 2) are shared in an SMT setting. The construction of the covert channel relies upon two actors: a sender and a receiver, executing on the same physical core (cf. Algo. 1 and Algo. 2 in Appendix). The receiver times the execution of a code sequence $\mathcal{C}$ ($\mathcal{C}$ has no possibility of any transient execution). In contrast, depending on whether the sender has to send bit $1$ or 00, it performs an additional execution involving transient execution. The presence or absence of transient execution causes contention in the concerned hardware assets, leading to a perceivable timing difference on the receiver side. Such a covert channel instantiated with SIMD-Vector transition transient executions has a bandwidth of $2$ KB/s with an accuracy of $87\%$.
@@ 7.3 Load Value Injection
In addition to the Leak attack type summarized in Tab. 2 and detailed in Sec. 7.1, the Floating Point operations relying upon an assist to carry out the denormalization operation forward incorrect data to subsequent transient instructions, before denormalization finishes. This scenario is a floating point flavor of classic load value injection (LVI) in transient executions [41]. However, unlike Sec. 7.1, the exact transient value injected into the transient operations is not architecturally visible. Hence, the adversary needs to mount LVI in two phases: ① offline phase where the adversary fuzzes the transient operation using different inputs and constructs the transient output, and ② online phase where the adversary uses the specially crafted inputs from ① that give a transient output of adversarial interest. On comparing the architectural result of the computation against the constructed LVI output, we observed both to be different, implying a successful transient value injection attack.
@@ 7.4 FP Assists due to FMA instructions
As stated in [41], denormalization requires a Floating Point assist (denoted by ASSISTS.FP in Tab. 1), which kicks in micro-operations to gradually underflow a floating point number close to 00. This causes a flush of the current in-flight micro-operations from the execution pipeline, thereby creating a transient window where non-denormalized floating-point values are forwarded to subsequent instructions speculatively, resulting in a transient window of executions on incorrect data, while the floating-point assist is denormalizing the concerned value. In [41], this transient window is used to construct a flavor of load-value injection exploits on the SSE2 instruction extension. Through Shesha however, we uncover existence of this transient execution path in other ISEs as well  [^3] Shesha also uncovers previously unreported combinations of AES and SSE instruction sequences (like AESDECLAST and MULSS sharing a Read-after-Write register dependency) to also incur Floating-Point assists., most notably in Fused-Multiply-Add (FMA) instructions like VFMADD132SD, VFMADD132PD, or exponentiation instructions like VSCALEFSS.
@@ 8 Exploiting Leakage from FMA
An interesting finding by Shesha is the transient nature of the Fused Multiply-Add (FMA) instruction set extension, which warrants a deeper investigation. Although the transient nature of AVX instructions is well-known [36, 41], the security implications of FMA (and its variants) instructions have not been studied in literature before. In this section, we develop on the findings of Sec. 7.4, and focus on the exploitability of previously unreported speculation in FMA execution.
@@ 8.2 Leaking arbitrary data from FMA instruction execution
From Sec. 7.4, we recall that Shesha has undiscovered the possibility of speculation in FMA execution. Such speculation, tied with the existence of buffers inside FMA execution engine, could leak to leakages. In this context, we test the following two hypotheses related to the transient behavior shown by the FMA instructions:
@@
With speculative execution in FMA execution engine (cf. Sec. 7.4), does FMA “Memory Access Units” transiently forward data to speculatively executing FMA instructions?
@@ 8.2.1 Intra-FMA Execution Engine Leakage
The attacker process forces speculation in FMA instruction (cf. Sec. 7.4) with memory operand. One example is VFMADD213PD zmm1, zmm2, zmmword PTR [rcx]), which multiplies the first and second operands and adds the value loaded by the third operand. Similar to the experiments detailed in Sec. 7.4, we force a floating-point assist by grooming zmm1 and zmm2 with carefully crafted denormal operands. Finally, we encode the entire width of zmm1 (which is the destination register) into a Flush+Reload covert channel to leak transiently forwarded data.
In our experiments, we use the primitives detailed in Tab. 4 to construct and enlarge the transient window to cover speculation in FMA. Overall, we do not observe the programmed victim memory leaking through the adversarial thread’s Flush+Reload channel. We attribute this absence of speculative forwarding of stale data to ① lack of speculative forwarding of stale data from MAUs to transiently executing FMA instructions, or ② the inability of constructed transient windows to cover transient execution of FMA instructions (which are inherently complex to execute). There is no evidence to favor one reason over another, though, and we leave this to future work.
@@ 8.2.2 FMA-AVX Execution Engine Leakage
To test hypothesis 2, we follow a similar setup as Sec. 8.2.1 with the exception that the adversarial thread engages the AVX execution unit. Interestingly, this time, we did observe statistically significant leakage. Listing 1 shows the Flush+Reload channel activity for the victim instruction VFMADD213PD zmm1, zmm2, zmmword PTR [rcx]. Note how the programmed victim memory (i.e. 0xfa) is clearly leaking to the adversary.
@@
To test the extent of such leakage, we divided the ISEs related to FMA in the following families and tested leakage for all instructions in each family: ① VFMADD* ($12$ instructions from $3$ sub-families: VFMADD132*, VFMADD213*, VFMADD231*), ② VFMADDSUB* ($6$ instructions from $3$ sub-families: VFMADDSUB132*, VFMADDSUB213*, VFMADDSUB231*), ③ VFMSUB* ($12$ instructions from $3$ sub-families (VFMSUB132*, VFMSUB213*, VFMSUB231*), ④ VFMSUBADD* ($6$ instructions from $3$ sub-families (like VFMSUBADD132*, VFMSUBADD213PD*, VFMSUBADD231PD*), ⑤ VFNMADD* ($12$ instructions from $3$ sub-families like VFNMADD132*, VFNMADD213*, VFNMADD231*), ⑥ VFNMSUB* (contains in-total $12$ instructions from $3$ sub-families like VFNMSUB132*, VFNMSUB213*, VFNMSUB231*), and ⑦ IFMA instructions (VPMADD52HUQ and VPMADD52LUQ) [^6] We present the comprehensive list in Listing 5 (Appendix D)..

### v2

@@ 4 Uncovered Transient Execution Paths
We ran a test campaign of $24$ hours, in which most observations were obtained within the first $8$ hours (refer to Algo. A in Appendix for description of the algorithm). Since we rely upon Intel Performance Monitoring Unit (PMU), which captures the events occurring in the system with certainty, the false positive rate of Shesha is zero. In this section, we detail the novel transient execution paths uncovered by Shesha as a result of the testing campaign. We also reverse engineer the root-cause of these transient execution paths and detail their effect on the overall behaviour of the system under test.
@@ 5 Shesha Design Choices
We now explain the choice of parameters by comparing the role of Particle behaviour and Swarm behaviour in Shesha. To do so, we define $6$ variants of Shesha wrt. hyperparameters $\beta$ (which controls Particle behaviour) and $\gamma$ (which controls Swarm behaviour). To compare between the variants, we define two metrics: ① time taken to find the first SIMD-Vector speculation, and ② time taken to find the first precision intermixing speculation. The reason for this (as detailed in subsequent sections) is because SIMD-Vector speculation is the easiest of the four (cf. Table 2) to discover. This is because of the large number of instructions under these ISEs, as well as no special requirement (like read-after-write dependencies, or specific floating-point arithmetic) needed to trigger this speculation. Precision intermixing, however, is on the other extreme, and is the hardest of the four to uncover. This is because (as we detail in Section 6.2) such speculation requires closely tied instructions that handle single/double precision respectively, and have read-after-write dependency.
Variants Shesha{β=1,γ=0} and Shesha{β=0,γ=1} are essentially equivalent to random sampling, with different sources. The former prioritizes mutating the chosen instruction within the particle itself by replacing it with a random instruction with probability $1$. The latter also works to similar goal, but replaces the chosen instruction with another instruction from other particle in the swarm. Then, variants Shesha{β=0.4,γ=0} and Shesha{β=0.1,γ=0} test the extent of particle behaviour by changing the probability $\beta$ of replacing a chosen instruction with a randomly sampled instruction. Finally, variants Shesha{β=0.1,γ=0.1} and Shesha{β=0.1,γ=0.4} test the extent of swarm behaviour by changing the probability $\gamma$ of replacing a chosen instruction with an instruction from another particle in the swarm. We summarize these in Table 3.
@@
As a result, in our final deployment of Shesha, we adopt a mixture of variant Shesha{β=0.4,γ=0} and variant Shesha{β=0.1,γ=0.4}. The first variant forms the Cognitive phase (cf. Section 3.4.2), while the second variant forms the Mixed phase (cf. Section 3.4.3). Finally, we note that variant Shesha{β=1,γ=0} simulates other fuzzing (essentially relying upon a random sampling strategy) based tools like [57, 40, 38, 41] had they supported ISEs like AVX/SSE/FMA and so on. For comparison however, we report the raw data from these works, albeit with different objectives over mostly x86 ISA. For instance, wrt. the speculative paths and ISA explored by [40], it is able to detect violations in about $5$ minutes (which is comparable to Shesha{β=0.4,γ=0}’s detection time for SIMD-Vector transient path detection). Likewise, in [38], the entire fuzzing campaign lasts about $20+$ CPU hours. On similar lines, the detection in [41] ranges from a minute to about an hour based on the type of speculation targeted 111[57] does not report convergence time.. One must note that the objectives of each of these tools (as presented in their corresponding papers) were different, and thus, they explore different subsets of ISAs. Shesha works with comparatively larger ISA set and thus explores wider range of instruction combinations and stimulates variety of specialized execution units and buffers.
@@ 6.1 SIMD-Vector Transition based SMC
SIMD and Vector instruction extensions like AVX, AVX2, AVX512EVEX, AVX512VEX, SSE, SSE2, SSE3, SSE4, SSE4a, and SSSE3 are central to several performance gains experienced in modern Intel processors. These ISEs allow greater than $64$ bit registers holding more than one data point, to be operated in parallel by the same instruction (hence the name SIMD type ISEs). To do so, modern IA-64 architectures define a register bank with sixteen $512$-bit wide registers. Interestingly, ISEs like AVX2/SSE allow differing representations of the same hardware register in two dimensions: ① width $w$ of the register, and ② the SIMD representation of the data contained in that width $w$. To elaborate on dimension ①, the same hardware registers can be interpreted as $512$-bit wide registers zmm0, zmm1, zmm2, ..., zmm15 (by ISEs like AVX512EVEX, AVX512VEX), as $256$-bit wide registers ymm0, ymm1, ymm2, ..., ymm15 (by ISEs like AVX2), and as $128$-bit wide regisers xmm0, xmm1, xmm2, ..., xmm15. Likewise, regarding dimension ②, the same data in any zmm/ymm/xmm register can be interpreted as differently packed integer/floating-point value. This discussion raises a concerning question: Given a processor allowing executions of ISEs requiring different representations of the same hardware, how are transitions between such ISEs handled? According to the Intel Software Developer Manual [24], the exact implementation to handle such transitions depends upon the exact micro-architecture under question. We defer the details of these implementations and our reverse-engineering approach to Appendix B. We make the following observation as the root cause of bad speculation (as in Table 2) encountered in modern Intel processors when transitioning from SIMD to Vector instructions (and vice-versa).
@@ 6.2 Precision based Transient Execution Paths
Precisely, for the discussed transient paths to kick in, the requirement is a set of at least two instructions with a Read-after-Write (RAW) dependency connecting them (using the xmm0 register in the above examples), without occurrence of an explicit conversion instruction (like CVTPD2PS or CVTPS2PD) in between. We note here that Write-after-Write (WAW) or Write-after-Read (WAR) dependencies do not trigger any of the following transient paths, simply because there is no need for any implicit precision conversion as the second instruction is writing to the dependent register and not reading from it. Writes from a single-precision instruction to a register holding a double-precision value are architecturally simply overwriting the registers, and thus incur no transient executions. We defer the detailed discussion on each of these transient execution paths and our reverse-engineering approach to Appendix C. We make the following observation as the root cause of all the bad speculation events when transitioning between single and double precision.
@@ 7.1 Transient Execution Window
Transient execution paths guarantee flushing of the in-flight micro-operations from the pipeline and re-steer the execution on the correct path. As such, the width of the transient window for any given transient execution path determines the number of micro-operations an attacker can issue on the transient path before the rollback of the instructions. Practically, a larger transient window allows for executing complex attack vectors in the transient path. Architecturally, no effect of any transient micro-operation is visible. Therefore, we need to rely upon micro-architectural effects to measure the transient window. In our case, we compute the transient window by counting the number of transient loads executed before the pipeline rollback. When no bad speculation (discovered by Shesha) occurs, we observe no activity in the transient path, implying absence of misprediction-based transient execution. Our results are summarized in Table 2. In comparison to the machine clears presented in [42], we conclude that the SIMD-Vector transition-based SMC exhibits a larger transient window than all other variants of transient executions through machine clears or assists (both uncovered here as well as in [42]). This observation directly ties in with the number of repetitions (for warming up the execution pipeline) an attacker needs to do on the transient path before the actual attack can be mounted (cf. Section 7.1). As evident, due to the largest transient window with minimum number of repetitions, SIMD-Vector transition-based SMC provides a very practical transient execution path (with effectively no repetitions required), thereby becoming an actual choice for an adversary to mount in-the-wild attacks.
Enlarging the transient window. As different transient execution paths have differing transient windows, not all of them would be directly suited to leak reliably. To circumvent this, a usual method is to warm up the pipeline before the actual attack; this can be done solely on the attacker’s end. This can be done by several repetitions (rep) of the same transient operations to maximize the transient window. To determine the extent of pipeline warm-up needed to leak data, we increase the repetitions of warm-up until we observe leakage from the Flush+Reload [59] covert channel. In Table 4, we provide a comparative analysis of the number of repetitions needed to warm up the pipeline. Lesser the number of required repetitions rep, more likely it is to find in-the-wild occurrence of the uncovered transient paths in in-the-wild code bases  [^2] We could not test the pipeline warmup and leakage rates for  [57] and [46] because of absence of Intel TSX on our test platforms..
@@ 7.2 SMT Covert Channel
We demonstrate a covert channel using contention in the affected hardware assets. Two processes are pinned to two logical cores of the same physical core. Hence, we assume Simultaneous Multi-Threading (SMT) or HyperThreading to be enabled for the covert channel construction to work since the affected assets identified (in Table 2) are shared in an SMT setting. The construction of the covert channel relies upon two actors: a sender and a receiver, executing on the same physical core (cf. Algo. 3 and Algo. 4 in Appendix). The receiver times the execution of a code sequence $\mathcal{C}$ ($\mathcal{C}$ has no possibility of any transient execution). In contrast, depending on whether the sender has to send bit $1$ or 00, it performs an additional execution involving transient execution. The presence or absence of transient execution causes contention in the concerned hardware assets, leading to a perceivable timing difference on the receiver side. Such a covert channel instantiated with SIMD-Vector transition transient executions has a bandwidth of $2$ KB/s with an accuracy of $87\%$.
@@ 7.3 Load Value Injection
In addition to the Leak attack type summarized in Table 2 and detailed in Section 7.1, the Floating Point operations relying upon an assist to carry out the denormalization operation forward incorrect data to subsequent transient instructions, before denormalization finishes. This scenario is a floating point flavor of classic load value injection (LVI) in transient executions [42]. However, unlike Section 7.1, the exact transient value injected into the transient operations is not architecturally visible. Hence, the adversary needs to mount LVI in two phases: ① offline phase where the adversary fuzzes the transient operation using different inputs and constructs the transient output, and ② online phase where the adversary uses the specially crafted inputs from ① that give a transient output of adversarial interest. On comparing the architectural result of the computation against the constructed LVI output, we observed both to be different, implying a successful transient value injection attack.
@@ 7.4 FP Assists due to FMA instructions
As stated in [42], denormalization requires a Floating Point assist (denoted by ASSISTS.FP in Table 1), which kicks in micro-operations to gradually underflow a floating point number close to 00. This causes a flush of the current in-flight micro-operations from the execution pipeline, thereby creating a transient window where non-denormalized floating-point values are forwarded to subsequent instructions speculatively, resulting in a transient window of executions on incorrect data, while the floating-point assist is denormalizing the concerned value. In [42], this transient window is used to construct a flavor of load-value injection exploits on the SSE2 instruction extension. Through Shesha however, we uncover existence of this transient execution path in other ISEs as well  [^3] Shesha also uncovers previously unreported combinations of AES and SSE instruction sequences (like AESDECLAST and MULSS sharing a Read-after-Write register dependency) to also incur Floating-Point assists., most notably in Fused-Multiply-Add (FMA) instructions like VFMADD132SD, VFMADD132PD, or exponentiation instructions like VSCALEFSS.
@@ 8 Exploiting Leakage from FMA
An interesting finding by Shesha is the transient nature of the Fused Multiply-Add (FMA) instruction set extension, which warrants a deeper investigation. Although the transient nature of AVX instructions is well-known [37, 42], the security implications of FMA (and its variants) instructions have not been studied in literature before. In this section, we develop on the findings of Section 7.4, and focus on the exploitability of previously unreported speculation in FMA execution.
@@ 8.2 Leaking arbitrary data from FMA instruction execution
From Section 7.4, we recall that Shesha has undiscovered the possibility of speculation in FMA execution. Such speculation, tied with the existence of buffers inside FMA execution engine, could leak to leakages. In this context, we test the following two hypotheses related to the transient behavior shown by the FMA instructions:
@@
With speculative execution in FMA execution engine (cf. Section 7.4), does FMA “Memory Access Units” transiently forward data to speculatively executing FMA instructions?
@@ 8.2.1 Intra-FMA Execution Engine Leakage
The attacker process forces speculation in FMA instruction (cf. Section 7.4) with memory operand. One example is VFMADD213PD zmm1, zmm2, zmmword PTR [rcx]), which multiplies the first and second operands and adds the value loaded by the third operand. Similar to the experiments detailed in Section 7.4, we force a floating-point assist by grooming zmm1 and zmm2 with carefully crafted denormal operands. Finally, we encode the entire width of zmm1 (which is the destination register) into a Flush+Reload covert channel to leak transiently forwarded data.
In our experiments, we use the primitives detailed in Table 4 to construct and enlarge the transient window to cover speculation in FMA. Overall, we do not observe the programmed victim memory leaking through the adversarial thread’s Flush+Reload channel. We attribute this absence of speculative forwarding of stale data to ① lack of speculative forwarding of stale data from MAUs to transiently executing FMA instructions, or ② the inability of constructed transient windows to cover transient execution of FMA instructions (which are inherently complex to execute). There is no evidence to favor one reason over another, though, and we leave this to future work.
@@ 8.2.2 FMA-AVX Execution Engine Leakage
To test hypothesis 2, we follow a similar setup as Section 8.2.1 with the exception that the adversarial thread engages the AVX execution unit. Interestingly, this time, we did observe statistically significant leakage. Listing 1 shows the Flush+Reload channel activity for the victim instruction VFMADD213PD zmm1, zmm2, zmmword PTR [rcx]. Note how the programmed victim memory (i.e. 0xfa) is clearly leaking to the adversary.
@@
To test the extent of such leakage, we divided the ISEs related to FMA in the following families and tested leakage for all instructions in each family: ① VFMADD* ($12$ instructions from $3$ sub-families: VFMADD132*, VFMADD213*, VFMADD231*), ② VFMADDSUB* ($6$ instructions from $3$ sub-families: VFMADDSUB132*, VFMADDSUB213*, VFMADDSUB231*), ③ VFMSUB* ($12$ instructions from $3$ sub-families (VFMSUB132*, VFMSUB213*, VFMSUB231*), ④ VFMSUBADD* ($6$ instructions from $3$ sub-families (like VFMSUBADD132*, VFMSUBADD213PD*, VFMSUBADD231PD*), ⑤ VFNMADD* ($12$ instructions from $3$ sub-families like VFNMADD132*, VFNMADD213*, VFNMADD231*), ⑥ VFNMSUB* (contains in-total $12$ instructions from $3$ sub-families like VFNMSUB132*, VFNMSUB213*, VFNMSUB231*), and ⑦ IFMA instructions (VPMADD52HUQ and VPMADD52LUQ) [^6] We present the comprehensive list in Listing 5 (Appendix E)..